# Quokka-based understanding of outflows (QED) - II. X-ray metallicity gradients as a signature of galactic wind metal loading

### v1

@@ 2.2.1 Generating photons using pyXSIM
Figure 2: A simulated Chandra/ACIS-I image in the 0.4 - 2 keV band of the snapshot shown in Figure 1; the colorbar shows photon count rate per pixel. We take $Z_{\mathrm{bg}}=0.5Z_{\odot}$ and use a 1000 Ms exposure time.
@@ 2.2.2 Generating mock observations using SOXS
Once the X-ray photons are generated, we use SOXS to produce mock images and spectra as they would be observed by the Chandra X-ray Observatory. SOXS uses a subset of these photons to project on the plane of the instrument and convolve with the instrument response to produce an events file. We use the response matrix (specified by Auxiliary Response File and Redistribution Matrix File) of the ACIS-I instrument which has a FoV of $16.9\times 16.9$ arcmin2 and a spectral resolution of 130 eV at 1.49 keV. The mock SOXS observation also needs an exposure time; we carry out our post-processing for exposure times of 1000, 100, 10, and 1 Ms in order to generate results with a range of SNRs. We summarise the parameters we use for the SOXS post-processing in Table 2.
Figure 2 shows a synthetic X-ray image for the same snapshot as Figure 1 as it would be observed by Chandra using the ACIS-I configuration. The axes represent the RA and Dec for a galaxy arbitrarily located at [$30^{\circ}$, $45^{\circ}$], while the color scale indicates photon counts per pixel integrated over energies from $0.4$ to $2.0$ keV, computed using the 1000 Ms results.
@@
For the purpose of generating spectra, we partition the simulation domain into 11 regions, mirroring the approach used by Lopez et al. (2020); we mark our regions in the bottom panel of Figure 1. The central region, labeled ‘D’, corresponds to the area where SNe are injected and extends up to a height of 489.03 pc (i.e., 256 cells), while the remaining 10 regions (‘A/B1’ to ‘A/B5’) cover the outflow zone, each spanning 733.55 pc (i.e., 384 cells) in height. We plot the spectra of the regions D and A1-5, again for the example case of the 110 Myr snapshot with $Z_{\mathrm{bg}}=0.5Z_{\odot}$, in Figure 3.
@@ 2.3 Fitting the X-ray spectrum
Due to the high-dimensional nature of the parameter space, we find that for certain values of the initial guess the fitter sometimes fails to find that true $\chi^{2}$ minimum; this problem is particularly likely to occur for regions far from the disc where the metallicity and SNR are lower. To mitigate this we carry out fits using a range of guesses for the initial elemental abundances from $Z/Z_{\mathrm{X,\odot}}=\max(Z_{\mathrm{bg}}-0.5,0)$ to $Z/Z_{\mathrm{X,\odot}}=Z_{\mathrm{bg}}+0.5$ in steps of 0.1, where $Z_{\mathrm{X},\odot}$ is the abundance of each species X at Solar metallicity. We then accept whichever fit returns the lowest $\chi^{2}$ value. Once we have these best-fitting values for each of the one- two- and three-plasma models, we accept as our final fit whichever of of these produces a reduced $\chi^{2}$ closest to unity for each region. For all the snapshots and values of $Z_{\mathrm{bg}}$ we consider in this paper, we find that for the central region ‘D’ the best fit is a three-plasma model, while all other regions are best fit by two-plasma models.
@@ 3 Results and Analysis
We are now in a position to discuss our main results. We use the procedures outlined in Section 2 to derive the metallicity and temperature values from the mock observations. As noted above, we carry out this analysis for $Z_{\mathrm{bg}}=0$, 0.2, 0.5, 1, and 2; we also analyse a control run in which we leave the density and temperature structure unchanged, but set the metallicity to a uniform value $Z=Z_{\odot}$ everywhere in the simulation domain. Except where otherwise noted we use the simulated observations for a 1000 Ms exposure so that we have essentially noise-free spectra.
@@ 3.1 Time variation
To ensure that our findings are not dependent on the specific simulation time we choose to analyse, we use the pipeline described in Section 2 to process four snapshots, corresponding to the state of the simulations at 100, 105, 110, and 115 Myr. We choose these specific times because the outflows have reached statistical steady-state after 100 Myr, and a 5 Myr interval between snapshots ensures that the results are not causally linked: the typical hot gas speed in the simulation is $\sim 1000$ km s-1, so 5 Myr interval is long enough hot gas generated at the midplane to exit the simulation domain at ±$\pm 4$ kpc.
@@ 3.2 Outflow metallicity patterns
The abundances of O, Ne, Mg, Si, and Fe plotted in the bottom panel of Figure 5, show similar patterns, which is expected given that we have assumed Solar-scaled abundances and thus the true abundance ratios are constant. This pattern is that the outflowing gas possesses significant metallicity gradients, with a peak in the central region D and a subsequent decline for A/B-1-5. Such a trend in metallicity is strikingly similar to that observed by Lopez et al. (2020) for S and Si in the wind of M82. We discuss the physical origin of this pattern and its implications in Section 4.
For comparison, we also show the “true” hot gas metallicity. There is some ambiguity about how to define this, since different choices of how to average over the metallicity distribution present in the real simulation data yield different outcomes. In order to produce something as close as possible to the quantity returned by the observations, we define our true metallicity gradient to be the X-ray luminosity-weighted mean metallicity. That is, we define the true metallicity as
@@ 3.3 Results for varying background metallicity
We can understand this behaviour as arising from a fundamental limitation of the commonly-used observational fitting procedure, which is that it assumes that the different temperature components share a common set of abundances. This assumption, while it is likely a practical necessity to avoid having the number of free parameters in the fit balloon out of control, is fundamentally incorrect, particularly near the disc, because the emitting medium here consists largely of bubbles of near-pure SN ejecta that are both very hot ($\sim 10^{7}$ K) and very metal-rich, and a slightly cooler (but still $\sim 10^{6}$ K) component of shocked ISM gas that is more metal-poor. Now let us consider the implications of this configuration for how the X-ray spectrum changes as we vary $Z_{\mathrm{bg}}$. The bubbles of SN-ejecta produce emission that is not tremendously sensitive to $Z_{\mathrm{bg}}$, because their metallicities are dominated by direct SN ejecta; they will be metal-rich even for $Z_{\mathrm{bg}}=0$. The shocked ISM component, on the other hand, has a luminosity that is highly sensitive to $Z_{\mathrm{bg}}$; low $Z_{\mathrm{bg}}$ makes this component dimmer compared to the emission from the fresh SN ejecta, while higher $Z_{\mathrm{bg}}$ makes it brighter. We can therefore understand the effect of $Z_{\mathrm{bg}}$ as changing the weight of the two components, each of very different intrinsic metallicity, that contribute to the overall spectrum.
@@
To confirm this hypothesis, we have experimented with generating spectra with emission from the most metal-rich gas, $Z>4Z_{\odot}$, artificially set to zero. We find that, when we do this, the anomalous behaviour that inferred abundances decrease with increasing $Z_{\mathrm{bg}}$ vanishes. Instead, consisent with our proposed hypothesis, we recover abundances that increase with $Z_{\mathrm{bg}}$ as expected.
@@
Finally, we note that neither the “true” X-ray-weighted metallicity nor the elemental abundances we retrieve from the X-ray spectrum fully agree with the mass-weighted metallicities computed directly from the simulation data, even if we limit ourselves to the hot phase. In particular, Figure 9 of QED I shows that for the hot phase, the mass-weighted metallicity is near-identical to the background metallicity for $Z_{\rm bg}\gtrsim 0.2$. Thus weighting by X-ray luminosity or the emission spectrum tends to accentuate the gradients compared to the mass-weighted average. For the true metallicity values, the averaging is biased towards regions of higher density, particularly because the luminosity is proportional to the square of the density. Therefore, even a small increase in emissivity due to increased background metallicity disproportionately biases the true metallicity average towards gas with a higher background, making the inferred metallicity appear higher than a mass-weighted value would suggest.
@@ 3.4 Impact of finite signal-to-noise ratio
Thus far we have restricted our analysis to spectra generated for an exposure time of 1000 Ms, so we are examining essentially noise-free data. It is therefore important to consider the more realistic case of observations with a finite signal-to-noise ratio (SNR), in order to determine what quality data are required to retrieve the essential features of the outflow metallicity pattern.
@@
Figure 8: Left: spectra for region ‘A5’ for the case of $Z_{\text{bg}}=0.5Z_{\odot}$ at 110 Myr measured at four different exposure times – 1000, 100, 10, and 1 Ms. The legend indicates the value of $\mathrm{SNR}_{\mathrm{obs}}$, the overall SNR we estimate from the exposure time (Equation 4). Right: Ne abundances as a function of height recovered from spectral fitting, for the same four SNRs shown in the left panel.
To generate results for a range of values of $\mathrm{SNR}_{\mathrm{obs}}$, we vary the exposure time from the initial 1000 Ms, decreasing it to 100, 10, and 1 Ms. For each of these cases we produce mock spectra with SOXS and fit using Sherpa, and we plot the resulting spectra for the A5 region and the spectrum-derived Ne abundance for all regions in the right and left panels of Figure 8, respectively; we again focus on Ne as an example, but the qualitative results are the same for other elements. In both the panels, the curves are labeled by $\mathrm{SNR}_{\mathrm{obs}}$. The blue curve ($\mathrm{SNR}_{\mathrm{obs}}=103.83$) is the case for a 1000 Ms exposure, and is identical to the red curve in Figure 7; exposure times of 100, 10, and 1 Ms yield $\mathrm{SNR}_{\mathrm{obs}}=32.84$, 10.50, and 3.43, respectively, as indicated in the legend. As one might expect, for poor SNRs, the deviation from the “perfect” spectrum is smallest for mid-energy bins ($E\sim 1$ keV) where Chandra is most sensitive, and increases rapidly for lower and higher energy bins. This may be the result of the relatively poor sensitivity of Chandra at these energies.
@@ 4.1 Negative metallicity gradients as evidence for metal loading
Our findings offer a clear explanation for the observation that galactic winds show $\alpha$ abundances that are enhanced relative to those of the ISM, but that this enhancement declines away from the galactic plane. This pattern occurs due to the distribution of metals among different gas phases and their respective contributions to X-ray emission. In the disc of the galaxy, recent supernova activity produces very hot gas that is highly enriched with $\alpha$ elements, leading to X-ray emission characteristic of elevated metallicities. At the same location there is a cold phase ($T<10^{4}$ K) of gas with much lower metallicity (but which nonetheless carries most of the total metal mass), this gas does not contribute to X-ray emission because of its low temperature, so the X-ray spectrum is dominated by the hot, very metal-rich gas from supernova ejecta. As the hot gas flows outward, however, it mixes with clouds of colder, metal-poorer gas from the surrounding ISM that have been entrained into the outflow. Over time and as the outflow moves away from the galactic plane, some of this cooler gas evaporates into the hot phase, diluting it and decreasing the mean metallicity of the X-ray emitting gas by an amount that increases with distance from the plane. This leads to a metallicity gradient where the X-ray emitting gas near the disc has higher metallicity than that further out in the wind. Although in our simulations we have assumed Solar-scaled abundances for simplicity, in reality, this trend should occur primarily for the $\alpha$ elements that are preferentially produced by type II SNe. This phenomenon explains the trends observed by Lopez et al. (2020).
@@ 4.2 Analytical Model of Phase Mixing
In this case, the metallicities of the cold gas vanish entirely, and we can directly read off the increase in the hot gas mass fraction due to evaporation from the ratio of the hot phase metallicities in the plane versus the wind. Thus for example, if observations of a galaxy reveal that the metallicity of the hot wind is half that of the hot midplane hot gas (i.e., $Z_{\text{h,w}}/Z_{\text{h,p}}=0.5$, roughly what Lopez et al. (2020) find for Si in the wind of M82), the simplified relation implies that the hot gas fraction in the midplane is approximately half that in the wind (i.e., $f_{\text{h,p}}/f_{\text{h,w}}\approx 0.5$), so the wind a few kpc from the galaxy consists roughly equally of gas that was heated at the midplane and escaped while hot, and cool gas that was entrained at the midplane but subsequently evaporated into the hot phase. However, we caution that this result will always be at least somewhat approximate, because we have seen that the elemental abundances inferred from X-ray measurements do not perfectly reflect the underlying abundances. For example, if we examine Figure 5, we can see that carrying out this exercise for O versus Ne would lead to quite different numerical values for $f_{\text{h,p}}/f_{\text{h,w}}$

### v2

@@ 2.2.2 Characteristics of the X-ray emitting gas
Figure 3: A simulated Chandra/ACIS-I image in the 0.4 - 2 keV band of the snapshot shown in Figure 1; the colorbar shows photon count rate per pixel. We take $Z_{\mathrm{bg}}=0.5Z_{\odot}$ and use a 10 Ms exposure time.
@@ 2.2.3 Generating mock observations using SOXS
Once the X-ray photons are generated, we use SOXS to produce mock images and spectra as they would be observed by the Chandra X-ray Observatory. SOXS uses a subset of these photons to project on the plane of the instrument and convolve with the instrument response to produce an events file. We use the response matrix (specified by Auxiliary Response File and Redistribution Matrix File) of the ACIS-I instrument which has a FoV of $16.9\times 16.9$ arcmin2 and a spectral resolution of 130 eV at 1.49 keV. The mock SOXS observation also needs an exposure time; we carry out our post-processing for exposure times of 10, 1, 0.1, and 0.01 Ms in order to generate results with a range of SNRs. We summarise the parameters we use for the SOXS post-processing in Table 2.
Figure 3 shows a synthetic X-ray image for the same snapshot as Figure 1 as it would be observed by Chandra using the ACIS-I configuration. The axes represent the RA and Dec for a galaxy arbitrarily located at [$30^{\circ}$, $45^{\circ}$], while the color scale indicates photon counts per pixel integrated over energies from $0.4$ to $2.0$ keV, computed using the 10 Ms results. We plot the spectra of the regions D and A1-5, again for the example case of the 110 Myr snapshot with $Z_{\mathrm{bg}}=0.5Z_{\odot}$, in Figure 4.
@@ 2.3 Fitting the X-ray spectrum
Due to the high-dimensional nature of the parameter space, we find that for certain values of the initial guess the fitter sometimes fails to find that true $\chi^{2}$ minimum; this problem is particularly likely to occur for regions far from the disc where the metallicity and SNR are lower. To mitigate this we carry out fits using a range of guesses for the initial elemental abundances from $Z/Z_{\mathrm{X,\odot}}=\max(Z_{\mathrm{bg}}-0.5,0)$ to $Z/Z_{\mathrm{X,\odot}}=Z_{\mathrm{bg}}+0.5$ in steps of 0.1, where $Z_{\mathrm{X},\odot}$ is the abundance of each species X at Solar metallicity. We then accept whichever fit returns the lowest $\chi^{2}$ value. Once we have these best-fitting values for each of the one-, two- and three-plasma models, we accept as our final fit whichever of these produces a reduced $\chi^{2}$ closest to unity for each region. For all the snapshots and values of $Z_{\mathrm{bg}}$ we consider in this paper, we find that for the central region ‘D’ the best fit is a three-plasma model, while all other regions are best fit by two-plasma models.
@@ 3 Results and Analysis
We are now in a position to discuss our main results. We use the procedures outlined in Section 2 to derive the metallicity and temperature values from the mock observations. As noted above, we carry out this analysis for $Z_{\mathrm{bg}}=0$, 0.2, 0.5, 1, and 2; we also analyse a control run in which we leave the density and temperature structure unchanged, but set the metallicity to a uniform value $Z=Z_{\odot}$ everywhere in the simulation domain. Except where otherwise noted we use the simulated observations for a 10 Ms exposure so that we have essentially noise-free spectra.
@@ 3.1 Time variation
To ensure that our findings are not dependent on the specific simulation time we choose to analyse, we use the pipeline described in Section 2 to process four snapshots, corresponding to the state of the simulations at 100, 105, 110, and 115 Myr. Our goal here is simply to assess the level of fluctuations in the spectra once the wind has reached a statistically steady state. We choose these specific times because the outflows have reached statistical quasi-steady-state after 100 Myr, and a 5 Myr interval between snapshots ensures that the results are not causally linked: the typical hot gas speed in the simulation is $\sim 1000$ km s-1, so 5 Myr interval is long enough hot gas generated at the midplane to exit the simulation domain at ±$\pm 4$ kpc.
@@ 3.2 Outflow metallicity patterns
The abundances of O, Ne, Mg, Si, and Fe plotted in the bottom panel of Figure 6 show similar patterns, which is expected given that we have assumed Solar-scaled abundances and thus the true abundance ratios are constant. This pattern is that the outflowing gas possesses significant metallicity gradients, with a peak in the central region D and a subsequent decline for A/B-1-5. Such a trend in metallicity is strikingly similar to that observed by Lopez et al. (2020) for S and Si in the wind of M82. We discuss the physical origin of this pattern and its implications in Section 4.
For comparison, we also show the “true” hot gas metallicity, reminding readers that, because our simulation assumes Solar-scaled abundances, the true (Solar-normalised) metallicity is the same for all elements. There is some ambiguity about how to define this, since different choices of how to average over the metallicity distribution present in the real simulation data yield different outcomes. In order to produce something as close as possible to the quantity returned by the observations, we define our true metallicity gradient to be the X-ray luminosity-weighted mean metallicity. That is, we define the true metallicity as
@@ 3.3 Results for varying background metallicity
We can understand this behaviour as arising from a fundamental limitation of the commonly-used observational fitting procedure, which is that it assumes that the different temperature components share a common set of abundances. This assumption, while it is likely a practical necessity to avoid having the number of free parameters in the fit balloon out of control, is fundamentally incorrect, particularly near the disc, because the emitting medium here consists largely of bubbles of near-pure SN ejecta that are both very hot ($\sim 10^{7}$ K) and very metal-rich, and a slightly cooler (but still $\sim 10^{6}$ K) component of shocked ISM gas that is more metal-poor. Now let us consider the implications of this configuration for how the X-ray spectrum changes as we vary $Z_{\mathrm{bg}}$. The bubbles of SN-ejecta produce emission that is not tremendously sensitive to $Z_{\mathrm{bg}}$, because their metallicities are dominated by direct SN ejecta; they will be metal-rich even for $Z_{\mathrm{bg}}=0$. The shocked ISM component, on the other hand, has a luminosity that is highly sensitive to $Z_{\mathrm{bg}}$; low $Z_{\mathrm{bg}}$ makes this component dimmer compared to the emission from the fresh SN ejecta, while higher $Z_{\mathrm{bg}}$ makes it brighter. We can therefore understand the effect of $Z_{\mathrm{bg}}$ as changing the weight of the two components, each of very different intrinsic metallicity, that contributes to the overall spectrum.
@@
To confirm this hypothesis, we have experimented with generating spectra with emission from the most metal-rich gas, $Z>4Z_{\odot}$, artificially set to zero. We find that, when we do this, the anomalous behaviour that inferred abundances decrease with increasing $Z_{\mathrm{bg}}$ vanishes. Instead, consistent with our proposed hypothesis, we recover abundances that increase with $Z_{\mathrm{bg}}$ as expected.
@@
Finally, we note that neither the “true” X-ray-weighted metallicity nor the elemental abundances we retrieve from the X-ray spectrum fully agree with the mass-weighted metallicities computed directly from the simulation data, even if we limit ourselves to the hot phase. In particular, Figure 9 of QED I shows that for the hot phase, the mass-weighted metallicity is near-identical to the background metallicity for $Z_{\rm bg}\gtrsim 0.2$. Thus weighting by X-ray luminosity or the emission spectrum tends to accentuate the gradients compared to the mass-weighted average. For the true metallicity values, the averaging is biased towards regions of higher density, particularly because the luminosity is proportional to the square of the density. Therefore, even a small increase in emissivity due to increased background metallicity disproportionately biases the true metallicity average towards gas with a higher background, making the inferred metallicity appear higher than a mass-weighted value would suggest. We explore this issue further in Section 4.3.
@@ 3.4 Impact of finite signal-to-noise ratio
Thus far we have restricted our analysis to spectra generated for an exposure time of 10 Ms, so we are examining essentially noise-free data. It is therefore important to consider the more realistic case of observations with a finite signal-to-noise ratio (SNR), in order to determine what quality data are required to retrieve the essential features of the outflow metallicity pattern.
@@
Figure 9: Left: spectra for region ‘A5’ for the case of $Z_{\text{bg}}=0.5Z_{\odot}$ at 110 Myr measured at four different exposure times – 10, 1, 0.1, and 0.01 Ms. The legend indicates the value of $\mathrm{SNR}_{\mathrm{obs}}$, the overall SNR we estimate from the exposure time (Equation 4). Right: Ne abundances as a function of height recovered from spectral fitting, for the same four SNRs shown in the left panel.
To generate results for a range of values of $\mathrm{SNR}_{\mathrm{obs}}$, we vary the exposure time from the initial 10 Ms, decreasing it to 1, 0.1, and 0.01 Ms. For each of these cases we produce mock spectra with SOXS and fit using Sherpa, and we plot the resulting spectra for the A5 region and the spectrum-derived Ne abundance for all regions in the right and left panels of Figure 9, respectively; we again focus on Ne as an example, but the qualitative results are the same for other elements. In both the panels, the curves are labeled by $\mathrm{SNR}_{\mathrm{obs}}$. The blue curve ($\mathrm{SNR}_{\mathrm{obs}}=103.83$) is the case for a 10 Ms exposure, and is identical to the red curve in Figure 8; exposure times of 1, 0.1, and 0.01 Ms yield $\mathrm{SNR}_{\mathrm{obs}}=32.84$, 10.50, and 3.43, respectively, as indicated in the legend. As one might expect, for poor SNRs, the deviation from the “perfect” spectrum is smallest for mid-energy bins ($E\sim 1$ keV) where Chandra is most sensitive, and increases rapidly for lower and higher energy bins. This may be the result of the relatively poor sensitivity of Chandra at these energies.
@@ 4.1 Negative metallicity gradients as evidence for metal loading
Our findings offer a clear explanation for the observation that galactic winds show $\alpha$ abundances that are enhanced relative to those of the ISM, but that this enhancement declines away from the galactic plane. This pattern occurs due to the distribution of metals among different gas phases and their respective contributions to X-ray emission. In the disc of the galaxy, recent supernova activity produces very hot gas that is highly enriched with $\alpha$ elements, leading to X-ray emission characteristic of elevated metallicities. At the same location there is a cold phase ($T<10^{4}$ K) of gas with much lower metallicity (but which nonetheless carries most of the total metal mass), this gas does not contribute to X-ray emission because of its low temperature, so the X-ray spectrum is dominated by the hot, very metal-rich gas from supernova ejecta. As the hot gas flows outward, however, it mixes with clouds of colder, metal-poorer gas from the surrounding ISM that have been entrained into the outflow. Over time and as the outflow moves away from the galactic plane, some of this cooler gas evaporates into the hot phase, diluting it and decreasing the mean metallicity of the X-ray emitting gas by an amount that increases with distance from the plane. This leads to a metallicity gradient where the X-ray emitting gas near the disc has higher metallicity than that further out in the wind. Although in our simulations we have assumed Solar-scaled abundances for simplicity, in reality, this trend should occur primarily for the $\alpha$ elements that are preferentially produced by core collapse SNe. This phenomenon explains the trends observed by Lopez et al. (2020).
@@ 4.2 Analytical Model of Phase Mixing
In this case, the metallicities of the cold gas vanish entirely, and we can directly read off the increase in the hot gas mass fraction due to evaporation from the ratio of the hot phase metallicities in the plane versus the wind. Thus for example, if observations of a galaxy reveal that the metallicity of the hot wind is half that of the hot midplane hot gas (i.e., $Z_{\text{h,w}}/Z_{\text{h,p}}=0.5$, roughly what Lopez et al. (2020) find for Si in the wind of M82), the simplified relation implies that the hot gas fraction in the midplane is approximately half that in the wind (i.e., $f_{\text{h,p}}/f_{\text{h,w}}\approx 0.5$), so the wind a few kpc from the galaxy consists roughly equally of gas that was heated at the midplane and escaped while hot, and cool gas that was entrained at the midplane but subsequently evaporated into the hot phase. However, we caution that this result will always be at least somewhat approximate, because we have seen that the elemental abundances inferred from X-ray measurements do not perfectly reflect the underlying abundances. For example, if we examine Figure 6, we can see that carrying out this exercise for O versus Ne would lead to quite different numerical values for $f_{\text{h,p}}/f_{\text{h,w}}$. We explore the reasons for these errors further in Section 4.3.
4.3 Why Do Biases in Derived Abundances Vary with Element?
Following our toy model of phase mixing, we now examine why X-ray derived abundances can differ systematically from the underlying metallicities, and do so differently for different elements. The key drivers of these differences are that (1) there is no single characteristic temperature for the “hot” ($T>10^{6}$ K) phase, but rather a broad distribution of temperature, (2) that temperature and metallicity are highly correlated, and (3) due to the different lines that contribute to the X-ray spectrum from $0.4-2$ keV, different elements are sensitive to gas at different temperatures, rather than simply reflecting some sort of mean X-ray weighted metallicity such as $Z_{\mathrm{met}}$ (Equation 1).
We illustrate the first two points using the case $Z_{\mathrm{bg}}=0.5Z_{\odot}$ as an example; for each region of this case we divide the gas hotter than $10^{6}$ K (the temperature range that drives the bulk of the X-ray emission; cf. Figure 2) into temperature bins 0.25 dex wide, and within each bin we compute the mass-weighted mean metallicity. We show the result for regions A1, A3, and A5 in Figure 11; other regions are qualitatively similar. A key point to take from these plots is that, except in region A5, “hot” gas spans more than an order of magnitude in temperature, and metallicity is highly-correlated with temperature, particularly close to the galactic plane. The origin of this correlation is simply that one of the most important cooling channels for the very hot gas produced by SNe, particularly at small $z$ when there has been little time for adiabatic expansion or radiative cooling, is dilution by mixing with cooler gas, and dilution simultaneously cools and lowers metallicity.
To illustrate how this interacts with the third point, regarding the different temperature sensitivity of different elements, we also in Figure 11 show the metallicities in each region inferred from the X-ray spectral fits for three sample elements, O, Ne, and Fe. We choose these three because, as shown in Figure 6, O and Ne show systematically the lowest and highest inferred abundances near the midplane, while Fe is intermediate between the two. We can understand the differences between these elements in light of Figure 11 as follows. First, note that the sensitivity of the spectrum to oxygen is largely through the prominent O viii feature at $\approx 0.65$ keV, and the strength of this feature is strongly influenced by the oxygen ionization balance, with the fraction of oxygen in the O viii reaching a maximum of $\approx 50\%$ at $T\approx 2.5\times 10^{6}$ K, and falling below 10% for $T\lesssim 1.2\times 10^{6}$ K or $T\gtrsim 6\times 10^{6}$ K (Sutherland & Dopita, 1993). Consequently, oxygen strongly reflects the abundances prevalent at the lower-temperature end of “hot gas”, $T\approx(2-3)\times 10^{6}$ K. By contrast, the greatest sensitivity to Ne is to Ne ix and Ne x features at $\approx 1$ keV, and the latter ionization state accounts for more than 10% of the available neon only at temperatures $T\approx(2.5-10)\times 10^{6}$ K, with a maximum of $\approx 50\%$ abundance at $T\approx 5\times 10^{6}$ K (Sutherland & Dopita, 1993). This makes Ne sensitive to significantly hotter gas than O, and so when abundant very hot gas is present – as Figure 11 shows it is in regions A1 and A3 – the metallicity inferred from Ne winds up reflecting the higher metallicity of this gas. By contrast iron is somewhat more complex, since the X-ray spectral feature most sensitive to iron is the complex of Fe L lines at $\approx 1$ keV, which can be produced by Fe in ionization states from Fe xviii to Fe xxiv. The lowest of these ionization states reaches $\gtrsim 10\%$ abundance at $T\approx 3\times 10^{6}$ K, leading to a behaviour that is somewhat intermediate between that of O and Ne, although the temperature dependence is more complex for Fe because the highest ionization state that produces Fe L emission remains abundant up to much higher temperatures, $T\approx 3\times 10^{7}$ K. The net effect of this complex dependence is that in the outer regions (e.g., A3 and A5) the oxygen abundance is closer to the X-ray weighted metallicity $Z_{\mathrm{met}}$, while near the disc (e.g., A1) the iron abundance is closer.
In summary, we can understand qualitatively why different elements are biased by different amounts based on the fact that they are sensitive to different ranges of gas temperature, and that gas at different temperatures really does have systematically different metallicities due to the effects of cooling by dilution. Thus, while metallicity declines with distance from the midplane, and all elements properly capture this trend, the discrepancy between the abundances derived for any particular element and the “true” metallicity value arises naturally from these complex combinations of atomic physics and the presence of a metallicity-temperature correlation in the gas. Given the complexity of this dependence, it seems unlikely that we will be able to acquire significantly more accurate results from spatially-unresolved spectra.
Figure 11: Mass-weighted metallicity in discrete temperature bins (purple curves), plotted together with horizontal lines showing the derived metallicities for oxygen (black dashed), neon (blue dashed), and iron (brown dashed). The red solid line shows the “true” X-ray luminosity-weighted metallicity $Z_{\mathrm{met}}$. We display three representative wind regions (‘A1’, ‘A3’, ‘A5’) in stacked panels. O vii and O viii lines are most sensitive to lower temperatures ($T\approx(2-3)\times 10^{6}$ K), while Ne and Fe trace hotter gas, which combined with the metallicity-temperature correlation lead metallicities derived based on the O abundance to be systematically smaller than those derived from Ne or Fe.